# RomanLens: The Role Of Latent Romanization In Multilinguality In LLMs

### v2

@@ Abstract
Large Language Models (LLMs) exhibit remarkable multilingual generalization despite being predominantly trained on English-centric corpora. A fundamental question arises: how do LLMs achieve such robust multilingual capabilities? We take the case of non-Roman script languages, we investigate the role of Romanization—the representation of non-Roman scripts using Roman characters—as a bridge in multilingual processing. Using mechanistic interpretability techniques, we analyze next-token generation and find that intermediate layers frequently represent target words in Romanized form before transitioning to native script, a phenomenon we term Latent Romanization. Further, through activation patching experiments, we demonstrate that LLMs encode semantic concepts similarly across native and Romanized scripts, suggesting a shared underlying representation. Additionally, for translation into non-Roman script languages, our findings reveal that when the target language is in Romanized form, its representations emerge earlier in the model’s layers compared to native script. These insights contribute to a deeper understanding of multilingual representation in LLMs and highlight the implicit role of Romanization in facilitating language transfer.
@@
Raj Dabre1,3,4,5
@@ 1 Introduction
The majority of modern Large Language Models (LLMs) Touvron et al. (2023); Dubey et al. (2024); Team et al. (2024) are trained predominantly on English-dominated corpora. Yet, they exhibit strong multilingual generalization across diverse languages Huang et al. (2023); Zhao et al. (2024a); Zhang et al. (2023). This raises a fundamental question: How do LLMs develop such robust multilingual capabilities despite their English-centric training?
To address this, we adopt a mechanistic interpretability perspective. Wendler et al. (2024) suggest that LLMs encode multilingual information within a shared, language-agnostic latent space, albeit with an inherent bias toward English due to training data composition and architectural choices. Further, Kojima et al. (2024) describe distinct phases in multilingual information processing: initial layers map language-specific lexical and syntactic representations to a language-independent semantic space, middle layers maintain this semantic abstraction, and final layers transform these representations into language-specific lexical and syntactic forms.
In this study, we focus on languages written in non-Roman scripts. We hypothesize that LLMs leverage romanized forms of non-Roman script languages as an intermediate bridge between their language-agnostic concept space and language-specific output representations. Romanization—the representation of non-Roman scripts using Roman characters—may facilitate this process by aligning non-English languages more closely with English. Supporting this, Jaavid et al. (2024) demonstrated that explicitly romanizing inputs improves model performance on multilingual tasks, suggesting an inherent alignment between romanized text and English representations. We investigate whether LLMs indeed use romanization as a bridge between language-agnostic concepts and language-specific outputs given its potential implications for understanding multilingual processing in LLMs.
@@ 2 Related Work
Recent studies have explored various aspects of LLMs’ multilingual behavior: examining whether English emerges as a latent language in English-centric LLMs Wendler et al. (2024), how the composition of training corpus mixtures influences latent representations Zhong et al. (2024) and how LLMs handle multilingual capabilitiesZhao et al. (2024b). Interpretability tools relevant to this work include logit lens Nostalgebraist (2020), tuned lens Belrose et al. (2023) and direct logit attribution Elhage et al. (2021) which are key tools for decoding intermediate token representations in transformer models. The logit lens applies the language modeling head to earlier layers without additional training, while the tuned lens in addition to this trains an affine mapping to align intermediate states with final token predictions. Direct logit attribution attributes logits to individual attention heads. This work focuses on the logit lens (Section 4.1) to investigate whether English-centric decoder only LLMs when prompted in a non-Roman language, processes via romanized latent states before producing native language text. Tuned lens is avoided as its training process might obscure the intermediate romanized states by aligning them to final native script outputs, potentially masking the phenomenon under investigation.
@@ 3 Background
We give a quick background of the transformer’s forward pass as well as the basics of mechanistic interpretability approaches such as logit lens and activation patching which we leverage in this paper.
@@
Decoder-only transformer models (Vaswani, 2017) employ a residual architecture to process input sequences through multiple layers, producing a sequence of hidden states (latents). These latents, whose dimensionality remains the same are updated iteratively across layers through transformer blocks , where indicates the layer index and is the final layer index. For next-token prediction, the final latent is transformed by an unembedding matrix to produce logit scores for vocabulary tokens which are then converted to probabilities via the softmax function (c.f. Appendix A).
@@ 3.2 Interpretability Tool: Activation Patching
In the context of activation patching, let denote the language of a word, denote the concept of a word and denote that word. Eg: If = cow and = ‘en’ then = ‘cow’. Similarly = ‘vache’. We use 5-shots translation prompts to create paired source and target prompt , with different concept, input language, and output language. Unless otherwise specified, and refer to the output languages of and , respectively.
@@ 4 Methodology
RQ2: How does the representation of semantic concepts in LLMs compare between native and romanized scripts of non-Roman languages? (Section 4.2)
@@
Prompt design. We design prompts that facilitate next-token () prediction from the given context (). This is adopted across all analysis setups. Appendix B provides a Hindi example of this prompt design, including its English translation and romanized form.
@@
Repetition task. We prompt the model to repeat a word in the same language given 5 in-context examples.
@@ 4.2 Patching With Romanized Representation Versus Native Representation
In all scenarios from the resulting next token distribution, we compute the probabilities i.e probability of source concept in target language, and i.e probability of target concept in target language. We track , i.e., the probability of the concept C occurring in language , by simply summing up the probabilities of all prefixes of and its synonyms in the next-token distribution (c.f. Appendix G). We analyze to evaluate how effectively a concept is encoded in a given source input language .
@@ 5 Experimental Settings
We adopt the dataset used in recent studies Dumas et al. (2024), extending it to include both native script and romanized versions of the languages considered in this study. Translations are performed using the Llama 3.3 70B model Dubey et al. (2024) and the translations are romanized using IndicXlit Madhani et al. (2023). All translations were manually validated to ensure data quality.
@@ 6 Results
We now share our observations about the role of Romanization in LLMs.
Native | Romanized | |
Single Source | ||
Multi Source |
@@ 6.1 Latent romanization
Among the three text completion tasks considered in the latent romanization experiment (Section 4.1), a relatively high latent romanization frequency is observed for the repetition task (Figure 6). The translation and cloze tasks function at a semantic level, whereas the repetition task is purely syntactic. This means that the repetition task being less complex, the model may reach a decision sooner, potentially in earlier layers and might express its prediction in romanized form in intermediate layers. This behavior could be attributed to language-specific neurons, responsible for native script processing, being predominantly concentrated in the initial and final few layers of LLMs Tang et al. (2024).
@@
This analysis reveals that LLMs encode semantic concepts similarly regardless of whether the input is in native or romanized script. Furthermore, this finding demonstrates that the model achieves comparable levels of language understanding when processing non-Roman languages in their romanized form as in their native script.
@@
We quantify the observations from Figure 2 by analyzing next-token predictions across layers using logit lens. In Figure 7, panels 7(c) and 7(d) illustrate that for translations into native scripts, target language tokens begin to emerge from layer 40 onward. Conversely, in panels 7(a) and 7(b), where the target language is in romanized script, target tokens appear 1–2 layers earlier. This pattern indicates that when processing non-Roman script languages, the model forms internal representations of target tokens in earlier layers for romanized script compared to native script. This trend is consistent across language pairs and models, as shown in Figures 16-18. This suggests that romanization facilitates faster progression toward language-specific embeddings.
@@ Discussion.
In our investigation of romanized representations in the latent layers, we conclusively identified romanized tokens in the last 6–7 layers of an LLM across various multilingual text completion tasks. Based on previous works in this field Wendler et al. (2024); Zhao et al. (2024b), in an English-centric decoder only LLM this region corresponds to the transition from an English-centric language-agnostic concept space to a language-specific space where the idea conceived in the concept space is expressed in the target language. Our findings suggest that romanization serves as a bridge between the concept space and the language-specific region for non-Roman script languages, an observation strongly supported by our analysis of six diverse writing systems. Romanization acting as a bridge could explain why romanization based script barrier breaking methods like Liu et al. (2024); Xhelili et al. (2024) work. Notably, we do not observe Latent Romanization in Chinese, likely due to its logographic script and relatively high-resource status.
@@ 7 Conclusion
Our findings show that LLMs implicitly use Romanization as a bridge for non-Roman scripts, exhibiting Latent Romanization in intermediate layers before switching to native scripts. Layerwise analyses reveal that semantic concepts are encoded similarly across native and Romanized inputs, indicating a shared internal representation. Moreover, when translating to a Romanized script, target words emerge earlier, highlighting Romanization as a structural link between language-agnostic concepts and language-specific output. While our study reveals initial insights into Latent Romanization, future work could focus on applying these findings to develop training strategies that enhance performance across diverse linguistic communities.
@@ 9 Ethics Statement
The code and datasets created in this work will be made available under permissible licenses. Generative AI systems were only used for assistance purely with the language of the paper, eg: paraphrasing, spell-check, polishing the author’s original content, and for writing boiler-plate code.
@@ Appendix G Computing Probabilities : Activation Patching Experiment
Cases of token overlap between i.e word representing source concept in target language and i.e word representing target concept in target language and their synonyms are excluded. Token overlap would cause ambiguity. Therefore in the final dataset,

### v3

@@ Abstract
Large Language Models (LLMs) exhibit strong multilingual performance despite being predominantly trained on English-centric corpora. This raises a fundamental question: How do LLMs achieve such multilingual capabilities? Focusing on languages written in non-Roman scripts, we investigate the role of Romanization—the representation of non-Roman scripts using Roman characters—as a potential bridge in multilingual processing. Using mechanistic interpretability techniques, we analyze next-token generation and find that intermediate layers frequently represent target words in Romanized form before transitioning to native script, a phenomenon we term Latent Romanization. Further, through activation patching experiments, we demonstrate that LLMs encode semantic concepts similarly across native and Romanized scripts, suggesting a shared underlying representation. Additionally, for translation into non-Roman script languages, our findings reveal that when the target language is in Romanized form, its representations emerge earlier in the model’s layers compared to native script. These insights contribute to a deeper understanding of multilingual representation in LLMs and highlight the implicit role of Romanization in facilitating language transfer. Code and data are available at https://github.com/AI4Bharat/Romanlens.
@@
Raj Dabre1,3,4,5444Work done during employment at NICT, Japan.
@@ 1 Introduction
The majority of modern Large Language Models (LLMs) Touvron et al. (2023); Dubey et al. (2024); Team et al. (2024) are trained predominantly on English-dominated corpora. Nonetheless, they exhibit strong multilingual performance across diverse languages Shi et al. (2023); Huang et al. (2023); Zhao et al. (2024a); Zhang et al. (2023). This raises a fundamental question: How do LLMs develop such robust multilingual capabilities despite their English-centric training?
To address this, prior work by Wendler et al. (2024) suggests that LLMs encode multilingual information within a shared, language-agnostic latent space, albeit with an inherent bias toward English due to training data composition and architectural choices. Building on this perspective, we investigate a complementary mechanism that may underlie multilingual processing, particularly for languages written in non-Roman scripts.
We hypothesize that LLMs leverage romanized forms of non-Roman script languages as an intermediate bridge between their language-agnostic concept space and language-specific output representations. Romanization—the representation of non-Roman scripts using Roman characters—may facilitate this process by aligning non-English languages more closely with English. Supporting this, Jaavid et al. (2024) demonstrated that explicitly romanizing inputs improves model performance on multilingual tasks, suggesting an inherent alignment between romanized text and English representations. We investigate whether LLMs indeed use romanization as a bridge between language-agnostic concepts and language-specific outputs given its potential implications for understanding multilingual processing in LLMs.
@@ 2 Related Work
Recent studies have explored various aspects of LLMs’ multilingual behavior: examining whether English emerges as a latent language in English-centric LLMs Wendler et al. (2024), how the composition of training corpus mixtures influences latent representations Zhong et al. (2024) and how LLMs handle multilingual capabilities Zhao et al. (2024b). Kojima et al. (2024) describe distinct phases in multilingual information processing: initial layers map language-specific lexical and syntactic representations to a language-independent semantic space, middle layers maintain this semantic abstraction, and final layers transform these representations into language-specific lexical and syntactic forms. Interpretability tools relevant to this work include logit lens Nostalgebraist (2020), tuned lens Belrose et al. (2023) and direct logit attribution Elhage et al. (2021) which are key tools for decoding intermediate token representations in transformer models. The logit lens applies the language modeling head to earlier layers without additional training, while the tuned lens in addition to this trains an affine mapping to align intermediate states with final token predictions. Direct logit attribution attributes logits to individual attention heads. This work focuses on the logit lens (Section 4.1) to investigate whether English-centric decoder only LLMs when prompted in a non-Roman script language, processes via romanized latent states before producing native language text. Tuned lens is avoided as its training process might obscure the intermediate romanized states by aligning them to final native script outputs, potentially masking the phenomenon under investigation.
@@ 3 Background
We give a quick background of the transformer’s forward pass, romanization and the basics of mechanistic interpretability approaches such as logit lens and activation patching which we leverage in this paper.
@@
Decoder-only transformer models (Vaswani, 2017) employ a residual architecture to process input sequences through multiple layers, producing a sequence of hidden states (latents). These latents, whose dimensionality remains the same, are updated iteratively across layers through transformer blocks , where indicates the layer index and is the final layer index. For next-token prediction, the final latent is transformed by an unembedding matrix to produce logit scores for vocabulary tokens which are then converted to probabilities via the softmax function (c.f. Appendix A).
3.1 Romanization
Transliteration is the conversion of text written in one script to another. Romanization is a subcategory of transliteration where the target script is English/Latin. Within romanization there are multiple romanization schemes available, each based on different considerations. One key aspect of romanization schema is if it is lossy or lossless. A lossless scheme is required in cases where we have to convert the output back to native script. Typically, deterministic transliterations are lossless, whereas natural transliterations are lossy.
Example:
The Hindi word for “flower” in Devanagari and its romanization:
[…]
Romanization: phool
@@ 3.3 Interpretability Tool: Activation Patching
In the context of activation patching, let denote the language of a word, denote the concept of a word and denote that word. For example, if = cow and = ‘en’, then = ‘cow’. Similarly = ‘vache’. We use 5-shot translation prompts to create paired source and target prompt , with different concept, input language, and output language. Unless otherwise specified, and refer to the output languages of and , respectively.
@@ 4 Methodology
RQ2: How does the representation of semantic concepts in LLMs compare between native and romanized scripts of non-Roman script languages? (Section 4.2)
@@
Prompt design. We design prompts that facilitate next-token () prediction from the given context (). This is adopted across all analysis setups. The prompts are designed around translation, repetition, and cloze tasks, as described below.
@@
Repetition task. We prompt the model to repeat a word in the same language given five in-context examples.
@@
These tasks cover a range of multilingual text completion setups. Among these, the repetition task is more syntactic in nature compared to translation and cloze tasks. Appendix B provides a Hindi example of prompts across each of the three tasks, along with their English translations and romanized forms.
@@ 4.2 Patching With Romanized Representation Versus Native Representation
In all scenarios from the resulting next token distribution, we compute the probabilities i.e. probability of source concept in target language, and i.e. probability of target concept in target language. We track , i.e., the probability of the concept C occurring in language , by simply summing up the probabilities of all prefixes of and its synonyms in the next-token distribution (c.f. Appendix G). We analyze to evaluate how effectively a concept is encoded in a given source input language .
@@ 5 Experimental Settings
Native | Romanized | |
Single Source | ||
Multi Source |
@@
We adopt the dataset used in recent studies Dumas et al. (2024), extending it to include both native script and romanized versions of the languages considered in this study. Translations are performed using the Llama 3.3 70B model Dubey et al. (2024) and the translations are romanized using IndicXlit Madhani et al. (2023), pypinyin Mozillazg (2024) and Unidecode Šolc (2025). All translations were manually validated to ensure data quality.
@@ 6.1 Latent romanization
Among the three text completion tasks considered in the latent romanization experiment (Section 4.1), a relatively high latent romanization frequency is observed for the repetition task (Figure 5). The translation and cloze tasks function at a semantic level, whereas the repetition task is purely syntactic. This means that the repetition task being less complex, the model may reach a decision of what to predict sooner, potentially in earlier layers and might express its prediction in romanized form in intermediate layers. This behavior could be attributed to language-specific neurons, responsible for native script processing, being predominantly concentrated in the initial and final few layers of LLMs, Tang et al. (2024) leaving the intermediate layers without them.
@@
This analysis reveals that LLMs encode semantic concepts similarly regardless of whether the input is in native or romanized script. Furthermore, this finding demonstrates that the model achieves comparable levels of language understanding when processing non-Roman script languages in their romanized form as in their native script.
@@
We quantify the observations from Figure 2 by analyzing next-token predictions across layers using logit lens. In Figure 7, panels 7(c) and 7(d) illustrate that for translations into native scripts, target language tokens begin to emerge from layer 40 onward. Conversely, in panels 7(a) and 7(b), where the target language is in romanized script, target tokens appear 1–2 layers earlier. This pattern indicates that when processing non-Roman script languages, the model forms internal representations of target tokens in earlier layers for romanized script compared to native script. This trend is consistent across language pairs and models (c.f. Figures 16-18 in the Appendix). This suggests that romanization facilitates faster progression toward language-specific embeddings.
@@ Discussion.
In our investigation of romanized representations in the latent layers, we conclusively identified romanized tokens in the last 6–7 layers of an LLM across various multilingual text completion tasks. Based on previous works in this field Wendler et al. (2024); Zhao et al. (2024b), in an English-centric decoder only LLM this region corresponds to the transition from an English-centric language-agnostic concept space to a language-specific space where the idea conceived in the concept space is expressed in the target language. Our findings suggest that romanization serves as a bridge between the concept space and the language-specific region for non-Roman script languages, an observation strongly supported by our analysis of six diverse writing systems. Romanization acting as a bridge could explain why romanization based script barrier breaking methods like Liu et al. (2024) and Xhelili et al. (2024) work. Notably, we do not observe Latent Romanization in Chinese, likely due to its logographic script and relatively high-resource status.
@@ 7 Conclusion
Our findings show that LLMs implicitly use Romanization as a bridge for non-Roman scripts, exhibiting Latent Romanization in intermediate layers before switching to native scripts. Layerwise analyses reveal that semantic concepts are encoded similarly across native and Romanized inputs, indicating a shared internal representation. Moreover, when translating into a Romanized script, target words emerge earlier, highlighting Romanization as a structural link between language-agnostic concepts and language-specific output. While our study reveals initial insights into Latent Romanization, future work could focus on applying these findings to develop training strategies that enhance performance across diverse linguistic communities.
@@ 9 Ethics Statement
The code and datasets created in this work will be made available under permissible licenses. Generative AI systems were only used for assistance purely with the language of the paper, e.g., paraphrasing, spell-check, polishing the author’s original content, and for writing boiler-plate code.
@@ Appendix G Computing Probabilities : Activation Patching Experiment
Cases of token overlap between i.e. word representing source concept in target language and i.e. word representing target concept in target language and their synonyms are excluded. Token overlap would cause ambiguity. Therefore in the final dataset,